Construct a truth table for each of these compound propositions. [6 marks]

a) p p b) p ⊕¬p

c) p ⊕¬q d) p ⊕¬q

e) (p q) (p ⊕¬q) f ) (p q) (p ⊕¬q)




1
Expert's answer
2021-11-22T05:44:08-0500

Let us construct the trush table for the following compound propositions:

a) "p \u2295 p"


"\begin{array}{||c||c|c|c|c|c||}
\hline\hline
p & p ⊕p \\
\hline\hline
0 & 0 \\
\hline
1 & 0\\

\hline\hline
\end{array}"


b) "p ⊕\neg p"


"\begin{array}{||c||c|c|c|c|c||}
\hline\hline
p & \neg p & p ⊕ \neg p \\
\hline\hline
0 & 1 & 1 \\
\hline
1 & 0 & 1\\

\hline\hline
\end{array}"


c) "p ⊕\neg q"


"\begin{array}{||c|c||c|c|c|c||}
\hline\hline
p & q & \neg 𝑞 & p ⊕\neg q \\
\hline\hline
0 & 0 & 1 & 1 \\
\hline
0 & 1 & 0 & 0\\
\hline
1 & 0 & 1 & 0\\
\hline
1 & 1 & 0 & 1\\
\hline\hline
\end{array}"


d) "\neg p ⊕\neg q"


"\begin{array}{||c|c||c|c|c|c||}
\hline\hline
p & q & \neg 𝑝& \neg q & \neg p ⊕\neg q\\
\hline\hline
0 & 0 & 1 & 1 & 0\\
\hline
0 & 1 & 1 & 0 & 1\\
\hline
1 & 0 & 0 & 1 & 1\\
\hline
1 & 1 & 0 & 0 & 0\\
\hline\hline
\end{array}"


e) "(p ⊕ q) ∨ (p ⊕\neg q)"


"\begin{array}{||c|c||c|c|c|c||}
\hline\hline
p & q & \neg q & p ⊕\neg q & p ⊕ q & (p ⊕ q) ∨ (p ⊕\neg q) \\
\hline\hline
0 & 0 & 1 & 1 & 0 & 1\\
\hline
0 & 1 & 0 & 0 & 1 & 1\\
\hline
1 & 0 & 1 & 0 & 1 & 1\\
\hline
1 & 1 & 0 & 1 & 0 &1\\
\hline\hline
\end{array}"


f ) "(p ⊕ q) \land (p ⊕\neg q)"


"\begin{array}{||c|c||c|c|c|c||}
\hline\hline
p & q & \neg q & p ⊕\neg q & p ⊕ q & (p ⊕ q) \land (p ⊕\neg q) \\
\hline\hline
0 & 0 & 1 & 1 & 0 & 0\\
\hline
0 & 1 & 0 & 0 & 1 & 0\\
\hline
1 & 0 & 1 & 0 & 1 & 0\\
\hline
1 & 1 & 0 & 1 & 0 &0\\
\hline\hline
\end{array}"